OpenStudy (anonymous):

What are the solutions of 2x2 + 16x + 34 = 0? Answer –8 ± 2i –8 ± i –4 ± 2i –4 ± i

OpenStudy (anonymous):

Quadratic formula.

OpenStudy (anonymous):

\[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

I still can't do it, Idk how to get the i

OpenStudy (anonymous):

\(i\) is just \(\sqrt{-1}\). Once you get the \(\sqrt{b^2-4ac}\) part solved, you get \(i\) by just substituting it for where \(\sqrt{-1}\) is.

OpenStudy (anonymous):

Quadratic problem is very easy, why ask the same question guys???

OpenStudy (anonymous):

@kr7210 , what is easy for you is not easy for others.

OpenStudy (anonymous):

Yeah....

OpenStudy (anonymous):

Keep in mind, kr, that you also struggle with things just as other people do. You struggle in the same way they do. All that is different is the subject.

OpenStudy (anonymous):

calm down @Limitless .... dont behave like a baby.... ;P... already show ur limits on chat!!!

OpenStudy (anonymous):

Kr, you are behaving immature by insulting others. You also got told to shut up in chat because other people found you immature and offensive. But this is not relevant to Katie's question.

OpenStudy (anonymous):

Katie, where are you stuck?

OpenStudy (anonymous):

Such a non-sense!!

OpenStudy (anonymous):

@kr7210 For your information, @Limitless is completely correct in this situation. People have the right to ask questions, and there shouldn't be an expectation that people will understand the way you, or anyone else, does. If you don't find these kind of problems interesting/challenging, I'm sure you can find another question more suited to your liking. In general, OpenStudy is an open, friendly environment. Please remember this in your interactions, thank you!

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

thanx for ur kind attention @cshalvey .. love to read this big stuff... i ll keep it in mind

OpenStudy (anonymous):

Cheers @kr7210 - much appreciated :)

OpenStudy (anonymous):

I can ask whatever I like which is why I'm on here.

OpenStudy (anonymous):

but dont ask whole ur assignment baby.... ;P ;D @Katiebae .. hahha

OpenStudy (anonymous):

thn why ask same question again nd again, u cannt get anything from previous same question? ;P

OpenStudy (anonymous):

It's timed, stop answering my questions if it's such a problem.

OpenStudy (anonymous):

lo... m gonna block u baby... dont worry...!! dont want to waste my tym on ur assignment... ;P ;D